OpenStudy (anonymous):

find any points of discontinuity for the two rational functions. describe any vertical or horizontal asymptotes and any holes. y= 2x^2+3 (over) x^2+2

OpenStudy (anonymous):

help @whpalmer4

OpenStudy (anonymous):

\[\frac{ 2x^2+3 }{ x^2+2 }\]

OpenStudy (anonymous):

Discontinuities will be at \(x^2+2=0\).

OpenStudy (anonymous):

Since \(x^2\) is positive for all real numbers, there are no discontinuities with real numbers.

OpenStudy (anonymous):

ok can you explain alittle more please

OpenStudy (anonymous):

Vertical asymptotes are caused by division by \(0\). The horizontal ones are found by letting \(x\) go to \(\pm \infty\)

OpenStudy (anonymous):

ok so why cant x^2 = -2

OpenStudy (anonymous):

Well, think about it this way. \(-\sqrt{2}\times -\sqrt{2} = 2\)

OpenStudy (anonymous):

You would have to let \(x =\pm\sqrt{2}i\)

OpenStudy (anonymous):

ok so for this problem there are no discontinuity

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it because we can not find a real number to make x^2+2 to equal 0

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

are there any holes or anything else in the graph for the rest of the problem??

OpenStudy (anonymous):

Polynomials are continuous. A polynomial divided by a polynomial is continuous so long as the bottom one is not zero.

OpenStudy (anonymous):

There may be horizontal asymptotes.

OpenStudy (anonymous):

ok how do we find that out?

OpenStudy (anonymous):

are you there???

OpenStudy (whpalmer4):

The horizontal asymptote we can find by dividing the first term of the numerator by the first term of the denominator. What do you get if you do that?

OpenStudy (anonymous):

I already told you. Find them by letting x go to positive or negative infinity.

OpenStudy (anonymous):

um not sure are we dividing 2x^2/x^2???

OpenStudy (whpalmer4):

when we let \(x\rightarrow\infty\) the first term is much bigger than all of the other terms, so it essentially reduces the fraction to \[\frac{2x^2}{x^2}\]What does that equal?

OpenStudy (anonymous):

umm 2???

OpenStudy (whpalmer4):

Yes.

OpenStudy (anonymous):

ok so the horizontal asymptote is 2??

OpenStudy (anonymous):

\[ \frac{ 2x^2+3 }{ x^2+2 } \cdot \frac{x^{-2}}{x^{-2}} = \frac{2+3x^{-2}}{1+2x^{-2}} \]

OpenStudy (whpalmer4):

Yes, \(y=2\) is the horizontal asymptote. @wio that's an elegant way of demonstrating it!

OpenStudy (anonymous):

ok so if we use wio equation it will get the same answer??

OpenStudy (whpalmer4):

Yes, his equation looks like this: \[\frac{2+\frac{3}{x^2}}{1+\frac{2}{x^2}}\] Hopefully you can see that as \(x\) gets large, the two little terms on the right get ever closer to 0.

OpenStudy (whpalmer4):

and if they get arbitrarily close to 0, then the fraction is arbitrarily close to 2/1.

OpenStudy (whpalmer4):

As you can see from this plot of @wio's fraction, even at only \(x=10\) the fraction is essential exactly 2.